Suppose the supply function in a market is and the current market price is p* = 4. What is the producer surplus?
b) Demand function is given by
i) Find the interval of prices for which demand is positive.
ii) Express total revenue TR = PQ as a function of price. When is total revenue maximized?
iii) For which price is the own-price elasticity equal to -1?
1
Expert's answer
2021-09-22T07:28:07-0400

Solution:

a.). Producer surplus = "\frac{1}{2} 
\times 
(B \times H)"

Qs = P2 = 42 = 16

Qd = 10 – P0.5 = 10 – 40.5 = 10 – 2 = 8

Producer surplus = "\frac{1}{2} 
\times 
(8 \times 16)" = 64

 

b. i.). The interval of prices for which demand is positive are 2 and 4

 

ii.). TR = P "\times" Q

Price = 4

To get quantity, derive the inverse demand function:

Q = 10 – P0.5

P = Q2 – 20Q + 100

TR = (Q2 – 20Q + 100) "\times" Q = Q3 – 20Q2 + 100Q

TR function = Q3 – 20Q2 + 100Q

 

Total revenue is maximized when the marginal revenue is equal to zero.

 

iii).  PED = "\frac{\triangle Q}{\triangle P} 
\times 
\frac{\ P}{\ Q}"

-1 = "1\times \frac{P}{8}"

P = -8

A price reduction of 8 is equal to the own-price elasticity of demand of -1